Aqueous sulfuric acid H2SO4 will react with solid sodium hydroxide NaOH to produce aqueous sodium sulfate Na2SO4 and liquid water H2O. Suppose 33.3 g of sulfuric acid is mixed with 12. g of sodium hydroxide. Calculate the minimum mass of sulfuric acid that could be left over by the chemical reaction. Round your answer to 2 significant digits.
Answer #1

H2504 + 2 NaOH = Naz S04 + 2H2O 03.30 33.39 = 0,3395 mol 4:980999/mo Moles of NagH = 39.997 g/mol A NaoH = - 12:00 - 0,3000m